# MotionMap: Representing Multimodality in Human Pose Forecasting

### v1

@@ Abstract
Human pose forecasting is inherently multimodal since multiple futures exist for an observed pose sequence. However, evaluating multimodality is challenging since the task is ill-posed. Therefore, we first propose an alternative paradigm to make the task well-posed. Next, while state-of-the-art methods predict multimodality, this requires oversampling a large volume of predictions. This raises key questions: (1) Can we capture multimodality by efficiently sampling a smaller number of predictions? (2) Subsequently, which of the predicted futures is more likely for an observed pose sequence? We address these questions with MotionMap, a simple yet effective heatmap based representation for multimodality. We extend heatmaps to represent a spatial distribution over the space of all possible motions, where different local maxima correspond to different forecasts for a given observation. MotionMap can capture a variable number of modes per observation and provide confidence measures for different modes. Further, MotionMap allows us to introduce the notion of uncertainty and controllability over the forecasted pose sequence. Finally, MotionMap captures rare modes that are non-trivial to evaluate yet critical for safety. We support our claims through multiple qualitative and quantitative experiments using popular 3D human pose datasets: Human3.6M and AMASS, highlighting the strengths and limitations of our proposed method. Project Page (link) Equal Contribution
@@ 1 Introduction
Figure 1: MotionMap uses heatmaps to depict a spatial distribution over the space of all possible motions. Local maxima imply that the corresponding motions have a higher likelihood of being a future motion for an observed pose sequence. MotionMap not only predicts a variable number of modes with the corresponding confidence, it explicitly encodes rare modes which could otherwise be averaged out.
Figure 2: We define a two stage pipeline for human pose forecasting. At first, we train a framework similar to an autoencoder to predict the ground truth and future motion (Sec: 4.3). However, at test time we do not have the future motion and its latent as input. Therefore, we train a heatmap model to predict MotionMap, which along with the codebook encodes the likely motions and their latents as a drop-in replacement (Sec: 4.4). At inference time, we use the predicted MotionMap to obtain latents corresponding to motions with a high likelihood and use it in tandem with the observed pose sequence to predict the future pose sequence (Sec: 4.5)
@@
We therefore begin by proposing an alternate paradigm to make human pose forecasting well-posed. Instead of attempting to learn an unbounded set of future motions, we encourage the pose forecasting model to explicitly learn future motions present in the observed data. Specifically, models should use the training set to learn different transitions from input to output pose sequences, and translate them to any unseen test sample. By doing so the problem is no longer ill-posed, but well-posed since for every input sequence there exists a fixed number of futures, bounded by the size of the training set. Moreover, explicitly learning these transitions from the training dataset allows us to identify unknown motions at test time.
@@
In this paper, we propose MotionMap, a novel approach that captures different future motions for each observation through a heatmap. Specifically, MotionMap interprets the heatmap as a spatial distribution over the space of all possible motion sequences in two dimensions. Different local maxima on this heatmap correspond to different possible future motions for an observed pose sequence (Fig: 1). This representation has the primary advantage of encoding and predicting a variable number of modes for each observed sequence. Furthermore, we show that this representation allows us to explicitly learn different transitions from the training distribution and translate them for unseen test samples. MotionMap allows us to incorporate two different measures of confidence/uncertainty in our pose forecasting framework. These are the confidence of each mode as well as the uncertainty in the prediction conditioned on the mode. Finally, by design, MotionMap is sample efficient in achieving mode coverage for an observed pose sequence.
We perform experiments on two popular human pose forecasting datasets: Human3.6M and AMASS. Specifically, we compare different methods for (1) sample efficiency, where we compare the metrics for a fixed number of predictions; and (2) the ability to recall transitions present in the observed data. We make two observations: (1) MotionMap has the highest sample efficiency across all methods, and (2) MotionMap can accurately recall transitions from the observed data for unseen test samples. We will make our code publicly available after peer review.
@@ 2 Related Work
Early approaches in human pose forecasting employed feed-forward networks [5, 6], and Recurrent Neural Networks (RNNs) to model the temporal aspects of the task [7, 8, 9]. Modeling temporal information implicitly is another effective method, commonly achieved by encoding each joint’s trajectory with the Discrete Cosine Transformation (DCT), which helps mitigate common failures of auto-regressive models [10]. Subsequent advancements incorporated Graph Convolutional Networks (GCNs) to better capture the spatial dependencies of human poses [11, 12, 13].Later, a specific GCN for this task was introduced that combined both temporal and spatial aspects within a single graph framework [14], and another GCN architecture in a two-stage prediction framework was presented, utilizing an initial guess network followed by a formal prediction network [15]. More recently, Transformers have shown effectiveness in capturing spatial and temporal dependencies, needed in human pose forecasting and researchers have incorporated them into their model designs in various configurations, including serial spatial and temporal attention blocks [16], parallel spatial and temporal blocks [17], and hybrid structures [18]. In addition, uncertainty in human pose forecasting has been studied [16], where they showed the impact of modeling homoscedastic uncertainty in the task. While all those approaches have yielded accurate short-term forecasts, they face significant challenges with relatively long-term predictions due to the inherent multimodality of human motion, where an observed sequence can lead to multiple futures.
@@
Multimodal pose forecasting acknowledges the inherent multimodality of human motion, aiming for a range of possible futures instead of a single outcome. This task has primarily been tackled through stochastic forecasting approaches  [19, 20, 21, 22, 23, 24, 25, 1, 26, 2, 27] using generative models such as Generative Adversarial Networks (GANs) [19, 20], Variational Auto-Encoders (VAEs) [21, 22, 23, 24], probabilistic latent variables [27] and more recently, Diffusion models [25, 1, 26, 2]. To promote diversity across the stochastic samples, DLow [3] introduced a new sampling strategy on top of conditional VAEs by generating multiple Gaussian distributions and then sampling latent codes from different Gaussian priors. GSPS [24] presented a two-stage sampling strategy, addressing the lower and upper body separately. DivSamp [4] proposed using the Gumbel-Softmax sampling strategy from an auxiliary space, and STARS [28] introduced an anchor-based sampling method. SLD [29] expanded this by projecting motion queries into a latent space to allow for more diverse motions, and MDN [30] proposed a transformer-based mechanism that enhances sampling diversity in the latent space. To generate more realistic forecasts, TCD [26] introduced a temporally-cascaded diffusion model that handles both perfect and imperfect observations, and BelFusion [1] presented a conditional latent diffusion model and recently, CoMusion [2] introduced a single-stage stochastic diffusion-based model using both Transformer and GCN architectures.
While these methods achieved notable results, they rely on stochasticity, learning an implicit distribution over the samples. As a result, these methods rely on large scale sampling for a wider mode coverage. Moreover, the number of samples to draw for different observations is unknown. In contrast, we attempt to determine the number of samples to draw per observation. We also propose a multimodal deterministic pose forecasting approach that explicitly captures all possible modes, ensuring diversity with fewer required samples. This method also enables us to quantify uncertainty and confidence in the predicted motions.
@@ 3 Multimodality in Human Pose Forecasting
We ponder upon the question: What is a mode in human pose forecasting? In the absence of a concrete definition, we define a mode as a set of motions corresponding to an action. For instance, walking could represent a class of motions and therefore a mode, and different variations of walking sequences could then be considered samples from this mode. Therefore, one could view multimodality as predicting a diverse set of actions that form a logical transition given an observed pose sequence.
Figure 4: Controllability. MotionMap can be used in combination with action labels for controllable pose forecasting. Since each pose sequence is associated with an embedding and action label, a spatial distribution over the space of motions is the same as over the action labels. This allows for the use of MotionMap to select of modes based on the confidence as well as user preference for the forecasted action. We illustrate this distribution over the space of motions $\leftrightarrow$ actions for an example input from the Human3.6M dataset.
However, this approach has two key limitations. First, there is no normalization of the skeletal size for different persons, implying that even if two pose sequences have similar motion, they would not be considered as each other’s multimodal ground truth if the skeletal sizes are very different. Second, the use of only the last frame from the input sequence may result in abrupt changes in the multimodal ground truth. This is because using one frame carries only positional information, i.e., no information about the motion in the input sequence. Therefore, we modify the definition of multimodal ground truth to (1) involve the last three frames, and (2) scale the skeletal dimensions of all other pose sequences to match that of the given pose sequence (Fig. 3).
We scale the skeleton by converting the pose from cartesian to spherical coordinates. Specifically, we use the spherical coordinate system to represent each joint (e.g. elbow) w.r.t its parent (e.g. shoulder). This allows us to maintain the angular positioning but modify the length of the child joint from its parent. By swapping this length with that of the reference skeleton, we obtain a new pose sequence the dimensions of which are scaled to match the reference skeleton. We provide the pseudo-code for scaling skeletal dimensions to match the target in the supplementary material (appendix: A).
@@ 4 Methodology
Now that we have defined our multimodal ground truths, let us discuss how we encode them through MotionMap, and use MotionMap in multimodal pose forecasting. We begin by describing the notation we use throughout the section.
4.1 Notation
Let $X$ be the sequence of observed poses consisting of $T_{o}$ frames $X=[x_{1},\dots,x_{T_{o}}]$. Each pose $x$ is of dimension $(J,3)$ where $J$ is the number of joints. Similarly, $Y$ is defined as the sequence of future poses with $T_{f}$ frames, $Y=[x_{T_{o}+1},\dots,x_{T_{o}+T_{f}}]$. We define the set of multimodal ground truths as $\bm{{Y}_{mm}}:\{{Y^{i}}|i:1,..,M\}$, where M{M} is the number of ground-truth modes. We note that the number of ground-truth modes varies for different samples. During inference, the pose forecaster receives $X$ as input and generates ${\bm{\hat{Y}_{mm}}}:\{\hat{Y^{i}}|i:1,..,\hat{M}\}$, where $\hat{M}$ is the predicted number of modes. We also note that M{M} and $\hat{M}$ could differ.
4.2 Overview
Our pose forecasting framework, depicted in Figure 2, consists of two trainable modules. We first have an autoencoder consisting of Gated Recurrent Unit encoders $\mathcal{E}_{\mathcal{X}}$, $\mathcal{E}_{\mathcal{Y}}$, a simple multilayer perceptron based uncertainty estimation module $\mathcal{U}$, and a GRUCell based decoder $\mathcal{D}$. We use their architectures from [1], and provide a detailed description in the supplementary material (appendix: B). The autoencoder is not only used for pose forecasting but also to obtain intermediate representations of different output pose sequences, which, as we shall see later, are used to learn the MotionMap. Our second trainable module is $\mathcal{H}$, which is used to predict the MotionMap for a given observation. Our training process therefore consists of two steps to learn the two modules.
@@ 4.3 Step 1: Autoencoder Module
The two encoders $\mathcal{E}_{X}$ and $\mathcal{E}_{Y}$ take as inputs $X$ and $Y_{i}$ respectively. If $X$ has multiple multimodal ground truths $\bm{{Y}_{mm}}$, then we select one at random. We expect that different ground truths are selected across different epochs. The outputs of these two networks are arrays $z_{x}$ and $z_{y}$ which are concatenated and passed through a simple two layer multilayer perceptron non-linearity giving us $f(z_{x}\bigoplus z_{y})$. Subsequently, this new array is passed to the decoder $\mathcal{D}$ to predict the entire sequence $\widehat{X\bigoplus Y_{i}}$. The reasoning behind predicting a concatenation of the input and multimodal ground-truth sequence is to avoid discontinuities and predict a smoother transition from the input pose sequence to the predicted pose sequence. This prediction along with the uncertainty predictions from $\mathcal{U}$ are optimized using the negative log-likelihood.
[…]
@@
We minimize this loss to jointly optimize the uncertainty and decoder, completing the training step for the autoencoder.
Figure 5: Ranking. Since MotionMap can predict variable number of modes with their associated confidences, our method also allows us to rank predictions. For instance, the highest ranked prediction (top row among $\hat{Y}$) closely matches the ground truth motion. However, rare modes (bottom row) are ranked low since the corresponding mode has a lower confidence.
@@ 4.4 Step 2: MotionMap Module
If we were to use the autoencoder as a pose forecaster at test time, we would be required to know the joint embedding $z_{x}\bigoplus z_{y}$, which is the input to the decoder. However, although we can compute $z_{x}$ since we know the input pose sequence, $z_{y}$ is nontrivial to obtain. While the literature utilizes various approaches stemming from generative models to obtain this latent, we propose MotionMap, which uses heatmaps and codebooks to represent various motion sequences. Intuitively, MotionMap can be interpreted as a learnt prior over the future motions for each sample, identifying the spread of modes and their associated likelihood. However, as the heatmap represents the spatial spread for two-dimensional coordinates, how can we trace these two dimensional coordinates back to $z_{y}$?
@@ 4.4.1 Dimensionality Reduction
We implement this by creating a mapping between the embedding $z_{y}$ and its two dimensional representation $h_{y}$ which is stored in a codebook $\mathcal{B}$. We obtain this two dimensional embedding by first encoding all future motions $Y$ using $\mathcal{E}_{y}$ to $z_{y}$. Next, we project all embeddings $z_{y}$ using t-SNE  [32] into two dimensions. Subsequently, we scale the two dimensional embeddings such that they span the entire size of the heatmap. Finally, we quantize the embeddings by rounding them to the nearest integer to obtain $h_{y}$. Since dimensionality reduction maps from the large volume source domain to a smaller volume target domain, there exist distinct $z_{y}$s that are mapped to the same $h_{y}$. Therefore, when creating the codebook, we map every $h_{y}$ to the mean of all $z_{y}$s that are mapped to it. To create the heatmap, we iterate over the list of all multimodal ground truths per sample, reduce each of them to their corresponding $h_{y}$s, and plot a Gaussian centered around it. Such an approach also ensures that very similar futures are mapped to the same $h_{y}$, avoiding duplicity of ground truths. As a result, rare modes are not suppressed in MotionMap.
Figure 6: Uncertainty. The Baro5 function used in [16] assumes (1) unimodality and a (2) temporally increasing trend in uncertainty. In contrast, we approach multimodality by decomposing the uncertainty into that of the mode and the forecast based on the mode (Sec. 5.3). We observe that this results in semantically richer uncertainty estimates. For instance, prediction 6 has a lower uncertainty in comparison to prediction 4 which involves fast changes in direction. Moreover, the pattern in uncertainty changes based on the motion. Finally, we predict the entire input-output sequence hence the model predicts lower uncertainty in reconstructing the input.

### v2

@@ Abstract
https://vita-epfl.github.io/MotionMap Equal Contribution
https://vita-epfl.github.io/MotionMap Equal Contribution
@@ 1 Introduction
Figure 1: MotionMap uses heatmaps to depict a spatial distribution over the space of motions. Local maxima imply that the corresponding motions have a higher likelihood of being a future motion for an observed pose sequence. MotionMap not only predicts a variable number of modes with the corresponding confidence, but it explicitly encodes rare modes that could otherwise be averaged out.
Figure 2: We define a two stage pipeline for human pose forecasting. At first, we train a framework similar to an autoencoder to predict the ground truth and future motion (Sec: 4.3). However, at test time we do not have the future motion and its latent as input. Therefore, we train a heatmap model to predict MotionMap, which along with the codebook encodes the likely motions and their latents as a drop-in replacement (Sec: 4.4). During fine-tuning and at inference time, we use the predicted MotionMap to obtain latents corresponding to motions with a high confidence and use it in tandem with the observed pose sequence to predict the future pose sequence (Sec: 4.5)
@@
We therefore begin by proposing an alternate paradigm to make human pose forecasting well-posed. Instead of attempting to learn an unbounded set of future motions, we encourage the pose forecasting model to explicitly learn future motions present in the observed data. Specifically, models should use the training set to learn different transitions from input to output pose sequences and translate them to any unseen test sample. By doing so the problem is no longer ill-posed, but well-posed since for every input sequence there exists a fixed number of futures, bounded by the size of the training set. Moreover, explicitly learning these transitions from the training dataset allows us to identify unknown motions at test time.
@@
In this paper, we propose MotionMap, a novel approach that captures different future motions for each observation through a heatmap. Specifically, MotionMap interprets the heatmap as a spatial distribution over the space of all possible motion sequences in two dimensions. Different local maxima on this heatmap correspond to different possible future motions for an observed pose sequence (Figure 1). This representation has the primary advantage of encoding and predicting a variable number of modes for each observed sequence. Furthermore, we show that this representation allows us to explicitly learn different transitions from the training distribution and translate them for unseen test samples. MotionMap allows us to incorporate two different measures of confidence/uncertainty in our pose forecasting framework. These are the confidence of each mode as well as the uncertainty in the prediction conditioned on the mode. Finally, by design, MotionMap is sample efficient in achieving mode coverage for an observed pose sequence.
We perform experiments on two popular human pose forecasting datasets: Human3.6M and AMASS. Specifically, we compare different methods for (1) sample efficiency, where we compare the metrics for a fixed number of predictions; and (2) the ability to recall transitions present in the observed data. We make two observations: (1) MotionMap has the highest sample efficiency across all methods, and (2) MotionMap can accurately recall transitions from the observed data for unseen test samples. We make our code publicly available on our project page.
@@ 2 Related Work
Architectures. Early approaches employed feed-forward networks [5, 6] and Recurrent Neural Networks (RNNs) to model the temporal aspects of the task [7, 8, 9]. Modeling temporal information implicitly is another effective method, commonly achieved by encoding each joint’s trajectory with the Discrete Cosine Transformation (DCT), which helps mitigate common failures of auto-regressive models [10]. Subsequent advancements incorporated Graph Convolutional Networks (GCNs) to better capture the spatial dependencies of human poses [11, 12, 13, 14, 15]. More recently, Transformers have shown effectiveness in capturing motion through serial spatial and temporal attention blocks [16], parallel spatial and temporal blocks [17], and hybrid structures [18].
@@
Multimodality. Recent pose forecasting methods acknowledge the inherent multimodality of human motion, aiming for a range of possible futures instead of a single outcome. This has been addressed through introducing stochasticity using generative models such as Generative Adversarial Networks (GANs) [19, 20], Variational Auto-Encoders (VAEs) [21, 22, 23, 24], probabilistic latent variables [25] and more recently, Diffusion models [26, 1, 27, 2]. To promote diversity across the samples, DLow [3] introduced a new sampling strategy on top of conditional VAEs by generating multiple Gaussian distributions and then sampling latent codes from different Gaussian priors. GSPS [24] presented a two-stage sampling strategy, addressing the lower and upper body separately. DivSamp [4] proposed using the Gumbel-Softmax sampling strategy from an auxiliary space, and STARS [28] introduced an anchor-based sampling method. Other methods such as SLD [29] projects motion queries into a latent space to allow for more diverse motions, and MDN [30] proposed a transformer-based mechanism that enhances sampling diversity in the latent space. To generate more realistic forecasts, TCD [27] introduced a temporally-cascaded diffusion model that handles both perfect and imperfect observations, and BelFusion [1] presented a conditional latent diffusion model and recently, CoMusion [2] introduced a single-stage stochastic diffusion-based model using both Transformer and GCN architectures.
Limitations. Although these methods achieve notable results, there exists certain limitations. First, diversity does not entail realism. While diversity based methods score high on diversity, the forecasts often lack coherence with the observation. Second, stochastic approaches commonly learn an implicit distribution over the multimodal future. As a result, they require large-scale sampling for a wider mode coverage. Moreover, the number of samples to draw for different observations is unknown. Further, are all samples equal, or are some more likely than the other? We argue for the need to have better representations for multimodality to address these limitations. However, we first begin by asking, what is multimodality in human pose forecasting?
@@ 3 Multimodality in Human Pose Forecasting
Without a concrete definition, we interpret a mode as a set of motions corresponding to an action which are likely for a given observation. For instance, walking represents a mode with various walking styles as samples. We add likely to avoid ambiguities, for example, related to left and right-handed actions within a mode. If the observation is left-handed, it is unlikely that the forecast should be right-handed. Therefore,
Multimodality in human pose forecasting refers to a diverse yet realistic set of future actions with a logical transition from an observed pose sequence.
@@
Figure 4: Controllability. MotionMap can also be used with auxiliary data such as action labels for controllable pose forecasting. Since each pose sequence is associated with an embedding and action label, a spatial distribution over the space of motions is the same as that over the action labels. This allows for the use of MotionMap to select modes based on the confidence as well as user preference for the forecasted action. We illustrate this distribution over the space of motions $\leftrightarrow$ actions for an example input from the Human3.6M dataset.
However, this approach has two main limitations. First, it does not normalize skeletal sizes across different individuals. Consequently, pose sequences with similar motions but different skeletal sizes are not recognized as multimodal ground truths for each other. Second, relying only on the final input frame can lead to abrupt transitions in multimodal ground truths, as a single frame contains positional but not motion-related information. To address these issues, we modify the multimodal ground truth definition by (1) incorporating the last three frames (2) scaling skeletons of other pose sequences to match the given sequence (Figure 3).
Skeleton scaling is achieved by converting poses from Cartesian to spherical coordinates. Specifically, we represent each joint (e.g., elbow) in spherical coordinates relative to its parent joint (e.g., shoulder), preserving angular positioning while allowing adjustments in joint length. By substituting the original joint lengths with those from the reference skeleton, we scale the entire pose sequence accordingly. Pseudo-code detailing this scaling procedure is provided in Appendix A.
@@ 4 Methodology
Using our definition of the multimodal ground truth, let us discuss how we encode them through MotionMap.
Notation. Let $X$ be the sequence of observed poses consisting of $T_{o}$ frames $X=[x_{1},\dots,x_{T_{o}}]$. Each pose $x_{i}$ is of dimension $(J,3)$ where $J$ is the number of joints. Similarly, $Y$ is defined as the sequence of future poses with $T_{f}$ frames, $Y=[x_{T_{o}+1},\dots,x_{T_{o}+T_{f}}]$. We define the set of multimodal ground truths as $\bm{{Y}_{mm}}:\{{Y^{i}}|i:1,..,M\}$, where ${M}$ is the number of ground-truth modes. We note that the number of ground-truth modes varies for different samples. During inference, the pose forecaster receives $X$ as input and generates ${\bm{\hat{Y}_{mm}}}:\{\hat{Y^{i}}|i:1,..,\widehat{M}\}$, where $\widehat{M}$ is the predicted number of modes. We also note that ${M}$ and $\widehat{M}$ could also differ.
Overview. Our pose forecasting framework, depicted in Figure 2, consists of two trainable modules. We have an autoencoder consisting of Gated Recurrent Unit encoders $\mathcal{E}_{\mathcal{X}}$, $\mathcal{E}_{\mathcal{Y}}$, a simple multilayer perceptron based uncertainty estimation module $\mathcal{U}$, and a GRUCell based decoder $\mathcal{D}$. We use their architectures from [1], and provide a detailed description in appendix B. The autoencoder is not only used for pose forecasting but also to obtain intermediate representations of different output pose sequences, which, as we shall see later, are used to learn the MotionMap. Our second trainable module is $\mathcal{H}$, which is used to predict the MotionMap for a given observation. Our training process therefore consists of two steps to learn the two modules.
@@ 4.1 Step 1: Autoencoder Module
The two encoders $\mathcal{E}_{X}$ and $\mathcal{E}_{Y}$ take as inputs $X$ and $Y$ respectively. If $X$ has multiple multimodal ground truths $\bm{{Y}_{mm}}$, then we select one $Y^{i}$ randomly, with different ground truths selected across different epochs. This is equivalent to a Monte-Carlo approximation of the multimodal distribution for each sample. The outputs of these two networks are arrays $z_{x}$ and $z_{y}$ which are concatenated and passed through a simple two layer multilayer perceptron non-linearity giving us $f(z_{x}\bigoplus z_{y})$. Subsequently, this new array is passed to the decoder $\mathcal{D}$ to predict the entire sequence $\widehat{X\bigoplus Y_{i}}$. The reasoning behind predicting a concatenation of the input and multimodal ground-truth sequence is to avoid discontinuities and predict a smoother transition from the input pose sequence to the predicted pose sequence. This prediction along with the uncertainty predictions from $\mathcal{U}$ are optimized using the negative log-likelihood.
Uncertainty. We use $f(z_{x}\bigoplus z_{y})$, the input to the decoder, to also predict the corresponding uncertainty of the prediction $\bm{\sigma}^{2}\in\mathbb{R}^{(T_{o}+T_{f})\times J}$. We use a normal distribution to model the uncertainty where both the mean and variance are predicted. If we define the mean square error per joint as $\text{error}=||\widehat{X\bigoplus Y_{i}}-X\bigoplus Y_{i}||_{2}^{2}$, the training objective is
@@
We minimize this loss to jointly optimize $\mathcal{U}$ and $\mathcal{D}$, completing the training step for the autoencoder. While we follow  [31] for simplicity, advanced techniques for heteroscedastic regression  [32, 33, 34, 35] are recommended.
Figure 5: Ranking. Since MotionMap can predict variable number of modes with their associated confidences, our method also allows us to rank predictions. For instance, the highest ranked prediction (top row among $\hat{Y}$) closely matches the ground truth motion. However, rare modes (bottom row) are ranked low since the corresponding mode has lower confidence.
@@ 4.2 Step 2: MotionMap Module
If we were to use the autoencoder as a pose forecaster at test time, we would be required to know the joint embedding $z_{x}\bigoplus z_{y}$, which is the input to the decoder. However, although we can compute $z_{x}$ since we know the input pose sequence, $z_{y}$ is nontrivial to obtain. While the literature utilizes various approaches stemming from generative models to obtain this latent, we propose MotionMap, which uses heatmaps and codebooks to represent various motion sequences. Intuitively, MotionMap can be interpreted as a learnt prior over the future motions for each sample, identifying the spread of modes and their associated likelihood. Given $\bm{\hat{Y}_{mm}}$, we encode each multimodal ground truth on a two-dimensional heatmap, allowing us to represent a variable number of futures per sample. However, how can we trace these future pose sequences to two dimensional coordinates and then back to $z_{y}$?
@@ 4.2.1 Dimensionality Reduction
We obtain two dimensional embeddings by first encoding all futures $Y$ across the dataset using $\mathcal{E}_{y}$ to $z_{y}$. Next, we project all embeddings $z_{y}$ using t-SNE  [36] into two dimensions. We choose t-SNE because it is a popular non-linear technique with an optimized implementation [37]. Subsequently, we scale the two dimensional embeddings such that they span the size of the heatmap. Finally, we quantize the embeddings by rounding them to the nearest integer to obtain $h_{y}$, of dimensionality two.
To create the heatmap, we iterate over each of the $M$ futures in $\bm{\hat{Y}_{mm}}$, reducing them to their corresponding $h_{y}$, and plot a Gaussian centered around it. Such an approach ensures that variable number of modes $M$ can be plotted per sample. Additionally, very similar futures are mapped to the same $h_{y}$, avoiding duplicity of ground truths. As a result, rare modes are not suppressed in MotionMap.
To go back to $z_{y}$, we create a mapping between the embedding $h_{y}$ and $z_{y}$ which is stored in a codebook $\mathcal{B}$. Since dimensionality reduction maps from the large volume source domain to a smaller volume target domain, there exist multiple $z_{y}$ that are mapped to the same $h_{y}$. Therefore, when creating the codebook, we map every $h_{y}$ to the mean of all $z_{y}$ that are mapped to it, giving us =$\mathcal{B}=\{h_{y}:\overline{z_{y}}\}$. The size of the codebook is independent of the dataset and depends on the level of quantization of the heatmap and the size of the embedding. For dimensionality $m\times m$, and using 32-bit floating point embeddings of length $n$, the size corresponds to $\approx 32m^{2}n$ bits. For $m=n=128$, the size is 64MB.
Figure 6: Uncertainty. Previous work [16] assumed homoscedasticity, learning generic uncertainty trends independent of the input. In contrast, we show that heteroscedastic modeling can result in semantically richer uncertainty estimates (Sec 5.3). We do this by decomposing uncertainty into that of the mode (through MotionMap) and the forecast conditioned on the mode (through the uncertainty module $\mathcal{U}$). For instance, prediction 6 has a lower uncertainty in comparison to prediction 4 because the latter involves fast changes in direction.